Size a square rapid-mixing basin, and provide specifications for mixer. The design flow rate is 7300 m3/d. The optimum G and Gt values are 750 s-1, and the Gt value of 20,000 – 30,000. The operating temperature is 10°C, the dynamic viscosity is 1.31 x 10-3 N∙s/m2, and the turbine shaft speed is 95 rpm with an impeller having 6 curved blades.

What will be the significance if you change your impeller design to a shrouded turbine with 6 curved blades?

What is the significance of the parameter G?
Answer #1

Step 1:

T=Gt/G

T=30000/750= 40sec

Q=7300 m3/day

Q=7300/86400=0.084 m3/sec

V=Qxt

V=0.084x40=3.36 m3

Step 2:

Consider H=1.25W

L=B=W

WXWX1.25W=3.36

1.25W3=3.36

W=1.39m

Consider it as 1.40m L=B=1.40M

D=1.25x1.4=1.75m

D.T=Volume/(Q)=1.4x1.4x1.75/0.084=40.83 sec

This design is satisfying the requirement

Step 3:

Now calculate the diameter of the impeller

P=G2V

P=750x750x1.31x10-3x1.4x1.4x1.75

P=2528Watts

N=95 rpm=95/60=1.58 rps

P=KtN3D5

2528=4.80x1000x1.583xD5

D=0.668m

Consider

Diameter of impeller=0.7m

Significance of parameter G is useful in determining how much the water is agitated in flash mixer, it is also useful in determining the power required and energy used